# TrustGeoGen: Scalable and Formal-Verified Data Engine for Trustworthy Multi-modal Geometric Problem Solving

### v1

@@ Abstract
Mathematical geometric problem solving (GPS) often requires effective integration of multimodal information and verifiable logical coherence. Despite the fast development of large language models in general problem solving, it remains unresolved regarding with both methodology and benchmarks, especially given the fact that exiting synthetic GPS benchmarks are often not self-verified and contain noise and self-contradicted information due to the illusion of LLMs. In this paper, we propose a scalable data engine called TrustGeoGen for problem generation, with formal verification to provide a principled benchmark, which we believe lays the foundation for the further development of methods for GPS. The engine synthesizes geometric data through four key innovations: 1) multimodal-aligned generation of diagrams, textual descriptions, and stepwise solutions; 2) formal verification ensuring rule-compliant reasoning paths; 3) a bootstrapping mechanism enabling complexity escalation via recursive state generation and 4) our devised GeoExplore series algorithms simultaneously produce multi-solution variants and self-reflective backtracking traces. By formal logical verification, TrustGeoGen produces GeoTrust-200K dataset with guaranteed modality integrity, along with GeoTrust-test testset. Experiments reveal the state-of-the-art models achieve only 49.17% accuracy on GeoTrust-test, demonstrating its evaluation stringency. Crucially, models trained on GeoTrust achieve OOD generalization on GeoQA, significantly reducing logical inconsistencies relative to pseudo-label annotated by OpenAI-o1. Our code is available at https://github.com/Alpha-Innovator/TrustGeoGen.
@@ 1 Introduction
Geometric Problem Solving (GPS), as a pivotal branch of mathematical reasoning or more broadly speaking, general reasoning, demands tripartite capabilities [40]: 1) extracting spatial relationships from visual diagrams, 2) performing mathematical reasoning in language modality, and 3) maintaining stepwise logical coherence where intermediate conclusions rigorously justify subsequent deductions. The development of trustworthy AGI systems fundamentally relies on reliably annotated multimodal data that not only bridges visual-textual modalities but also preserves closed-loop reasoning chains. This requirement stems from a critical axiom: a geometrically valid solution demands both correct answers and air-tight logical pathways, as human experts reject proofs with leap-of-faith***Here it means an unjustified or unsubstantiated jump in reasoning, where a conclusion is reached with insufficient logical steps or evidence. arguments even when final results are accurate.
@@
The recent advancement of Multimodal Large Language Models (MLLMs) [30, 22, 28, 4, 24, 14, 8, 20, 40, 27, 37, 11, 45, 12] has revealed promising capabilities in geometric reasoning, demonstrating competence in solving elementary geometric problems. Parallel developments in external formalized expert systems [7, 18, 44, 21] have extended problem solving capabilities to more complex geometric configurations, with highly specialized models [35, 9, 43] even approaching challenges of the International Mathematical Olympiad (IMO) level. However, these achievements are critically dependent on high-quality geometric data, currently in short supply. The existing datasets are mainly derived from manual annotations of secondary school textbook problems [6, 7, 15, 44, 21], and this process is limited in scale and diversity. Although recent integration of MLLMs (e.g. GPT series [2, 23, 24]) for synthetic question-answer generation [46, 14] has partially ameliorated data scarcity concerns, significant challenges persist in ensuring stepwise solution validitySpecifically, automatically generated procedural labels frequently contain latent inconsistencies, such as misapplied geometric axioms or diagram-coordinate mismatches., which may undermine the logical coherence of reasoning chains during model training.
However, current geometric datasets exhibit four deficiencies: 1) Modality Fragmentation: existing datasets suffer from incomplete modality alignment, lacking comprehensive annotations that synchronize images, textual captions, questions, and chain-of-thought reasoning processes within unified instances. 2) Scalability Issues: Most approaches rely on repurposing human-authored problems from textbooks or examinations rather than generating novel, curriculum-aligned challenges. 3) Cost Constraints: Manual annotation requires costly geometry expertise, while generating solutions through MLLMs consumes excessive computational resources. 4) Trust Deficits: Model-generated solutions lack reliable automated verification mechanisms, introducing risks of undetected logical errors or misaligned reasoning steps.
To this end, we develop TrustGeoGen, a formal language-verified geometric data generation engine that autonomously produces geometric problems, diagrams, and stepwise solutions through four integrated components: 1) a Constructor to generate constraint-satisfying problem premises and diagrams, 2) a Reasoner expanding geometrically valid reasoning graphs with rigorous node verification, 3) a Sampler leveraging the algorithm GeoExplore to extract high-quality reasoning paths, which are later converted into problems by pairing initial premises with derived conclusions, and 4) a Translator rendering formal specifications into natural language. Furthermore, we improve the problem complexity through a bootstrapping mechanism in TrustGeoGen, where formally verified geometric premises in prior stages are recursively fed back to the Constructor to iteratively append new premises. This closed-loop process enables systematic escalation of difficulties while maintaining logical consistency. Additionally, we refine the algorithm GeoExplore to simultaneously explore alternative solution paths and self-correcting reasoning trajectories, generating problems with multiple valid solutions and self-reflective traceback data.
Our proposed TrustGeoGen ensures: 1) Modality Completeness with synchronized dual-form (formal & informal) captions, questions, solutions, and diagrams, 2) Scalability in generating geometry problems by bootstrapping mechanisms without reliance on existing data source and 3) Verifiability through formal validation of all solution steps in its integrated reasoner. Based on TrustGeoGen, we develop a dataset GeoTrust containing 200K samples and a compact testset GeoTrust-test with 240 samples. Experimental results demonstrate state-of-the-art reasoning models DeepSeek-R1 [10] and OpenAI-o1 [25] achieve only 45.00% and 49.17% accuracy respectively on GeoTrust-test. Furthermore, training with our strictly verified reasoning step annotation data proves to be more effective compared to using pseudo-labelsHere, pseudo-label means that the solution to a given question is annotated using SOTA MLLMs (e.g. OpenAI-o1 [25]). with unverifiable reasoning processes. Additionally, our training data from GeoTrust enhances model performance on the out-of-distribution (OOD) testset GeoQA [6], highlighting its broader generalization capability. In a nutshell, our contributions are as follows.
@@
  • We develop TrustGeoGen, a formalized geometric data generation engine, producing scalable and multimodal-integrated data (geometric diagrams, formal & natural language captions, questions and solutions) with trustworthy reasoning guarantees via formal verification.
  • By introducing a bootstrap mechanism, we iteratively generate data with more complex premises and longer reasoning chains, enabling automated complexity scaling.
  • Our GeoExplore series algorithms generates multi-solution problems and self-reflective traceback data, enriching diversity while ensuring trustworthiness.
  • We construct the dataset GeoTrust with 200K samples and GenTrust-test testset, dynamically generated without dependency on prior data sources, ensuring unbiased and extensible evaluating capabilities.
  • Experiments demonstrate that existing MLLMs significantly underperform on high-difficulty geometric reasoning tasks, and our verified data exhibit effectiveness and generalization as training source.
@@ 2.1 Data Engine
In the process of generating multi-modal geometric reasoning data, it employs a rule-based data construction method augmented by a geometric verifier to synthesize high-quality and scalable datasets. This approach ensures that the data is accurate, controllable, and scalable. As shown in Fig. 1, our framework consists of four components: Constructor, Reasoner, Sampler, and Translator.
Constructor starts the data construction process by building geometric scenes. It incrementally adds new conditions to a basic geometric scene, thus constructing increasingly complex scenes. In TrustGeoGen, a basic geometric scene consists of simple geometric shapes (e.g., triangles, circles, parallelograms) and their associated numerical information (e.g., side lengths, radii, angles). When a base scene is created, it contains an initial set of geometric premises $P_{base}=\{p_{i}^{g},p_{j}^{d}\}$, where $p_{i}^{g}$ represents geometric relationship conditions (e.g., $AB//CD$) and $p_{j}^{d}$ represents numerical conditions (e.g., $\angle BAC=30^{\circ}$). Subsequently, the Constructor incrementally adds new premises to the base scene, increasing its complexity while expanding the premise set $P$. Notably, the Constructor requires that each newly added premise must be connected to the preceding scene to ensure the overall geometric scene remains cohesive and connected. By controlling the number of new premises added, the Constructor generates geometric scenes of varying complexity levels. Each constructed scene is then validated by a geometric compiler [31] to ensure the validity of its geometric relationships. Finally, the process yields a geometric scene with a specified level of complexity and an expanded premise set $P=\{p_{i+m}^{g},p_{j+n}^{d}\}$.
Reasoner leverages a predefined set of geometric theorems to infer new states from premises generated by the constructor, formulating a reasoning graph:
$\small\mathcal{G}=(S,S_{0},R,\hookrightarrow)$ (1)
  • $S$ is the set of states, where each state $s\in S$ represents a derived conclusion or fact in the reasoning process.
  • $S_{0}=P\subseteq S$ is the set of initial states (the premises $P$ generated by the constructor).
  • $R$ is the set of rules, where each rule $r\in R$ defines a logical inference relationship between states.
[…]
Sampler operates on the reasoning graph $\mathcal{G}$ by sampling a target state $s_{n}$ and subsequently searching for its path:
$\small\mathcal{P}=\{(S_{i-1},r_{s},s)\mid\forall s\in S_{i},S_{i-1}\xhookrightarrow{r_{s}}s,i=n,\dots,1\}$ (3)
where each triplet $(S_{i},r_{s},s)$ denotes a specific transition during the reasoning process, derived by applying rule $r_{s}$ to state set $S_{i-1}$ to produce state $s$. Any state $s_{n}$ within the state set $S$ can be treated as a problem to solve in order to derive a reasoning path $\mathcal{P}$. The sampler is initialized with a state $s\in S_{n}$ to retrieve the corresponding transition rule $r_{s}$ and the upstream dependent states $S_{i}$. This process is repeated iteratively until the state set $S_{i}$ consists entirely of initial premises $S_{0}$. At that point, the full reasoning path has been successfully constructed. Notably, during the graph construction process in the Reasoner, once a state $s$ is derived through rule $r_{s}$, no additional state transition relations are appended to $s$. Consequently, the Sampler is always able to identify a unique inference rule $r_{s}$ corresponding to any state $s$, thereby guaranteeing the absence of cycles in the reasoning path.
To ensure the quality of the reasoning paths, the Sampler employs two metrics to filter out unsuitable paths. The first metric evaluates the length of the reasoning path $\mathcal{P}$ against a predefined threshold $\tau_{l}$. If the length of $\mathcal{P}$ exceeds $\tau_{l}$, the path is considered difficult enough. Formally, this condition is expressed as $|\mathcal{P}|\geq\tau_{l}$, where $|\mathcal{P}|$ denotes the number of transitions in the reasoning path $\mathcal{P}$. The second metric assesses the ratio of the states in the reasoning path $\mathcal{P}$ that are derived from the initial premises set $S_{0}$. If this ratio exceeds a predefined threshold $\tau_{r}$, the path is deemed sufficient enough. This condition is formally defined as $\frac{|S_{P}|}{|S_{0}|}\geq\tau_{r}$, where $|S_{P}|$ represents the number of initial premises used in the reasoning path $\mathcal{P}$, and $|S_{0}|$ is the number of all initial premises. By applying these two metrics, the Sampler ensures that only high-quality reasoning paths are retained, thereby maintaining the integrity and usefulness of the generated data. The detailed procedure of the aforementioned reasoning path exploration and filtering algorithm, GeoExplore, is presented in Algorithm 1.

### v2

# TrustGeoGen: Formal-Verified Data Engine for Trustworthy Multi-modal Geometric Problem Solving

@@ Abstract
Mathematical geometric problem solving (GPS) demands verifiable logical coherence and multimodal reasoning capabilities. While large language models (LLMs) have shown rapid progress in GPS, their advancement is hindered by the lack of reliable benchmarks and systematic methodologies. A critical challenge is the inherent hallucination in LLMs, which leads to synthetic GPS datasets that are often noisy, unverified, and self-contradictory. To address this, we introduce TrustGeoGen, a data engine that generates formally verified geometric problems to establish a principled and trustworthy benchmark. Our engine integrates four key innovations: 1) Multimodal Alignment, which synchronizes the generation of diagrams, text, and step-by-step solutions; 2) Formal Verification, ensuring all reasoning paths are rule-compliant; 3) Connection Thinking, bridging formal deduction with human-like logical steps; and 4) our GeoExplore series algorithms, which produce diverse problem variants with multiple solutions and self-reflective backtracking. Using this engine, we create the GeoTrust-200K dataset and the corresponding GeoTrust-test benchmark, both with guaranteed cross-modal integrity. Experiments reveal that state-of-the-art models achieve only 45.83% accuracy on GeoTrust-test, highlighting its significant challenge. Furthermore, training on our synthesized data substantially improves model performance on GPS tasks, with strong generalization to out-of-domain (OOD) benchmarks. Our code and data are available at https://github.com/Alpha-Innovator/TrustGeoGen.
Index Terms:
Data Engine, Geometry, Multi-modal
@@ 1 Introduction
Geometric problem solving (GPS) geox/Xia2024GeoXGP , as a pivotal branch of mathematical reasoning or more broadly speaking, general reasoning, demands tripartite capabilities: 1) extracting spatial relationships from visual diagrams, 2) performing mathematical reasoning in language modality, and 3) maintaining stepwise logical coherence where intermediate conclusions rigorously justify subsequent deductions. The development of trustworthy AGI systems fundamentally relies on reliably annotated multimodal data that not only bridges visual-textual modalities but also preserves closed-loop reasoning chains. This requirement stems from a critical axiom: a geometrically valid solution demands both correct answers and air-tight logical pathways, as human experts reject proofs with leap-of-faith [^1] Here it means an unjustified or unsubstantiated jump in reasoning, where a conclusion is reached with insufficient logical steps or evidence. arguments even when final results are accurate.
The recent advancement of multimodal large language models (MLLMs) (math-llava/shi2024math, ; mathvista/lumathvista, ; peng2024chimera, ; qwen-vl/bai2023qwen, ; openai2024gpt4o, ; g-llava/gao2023g, ; chen2024far, ; lu2024deepseek, ; geox/Xia2024GeoXGP, ; wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels, ; internlmxcomposer2, ; internlmxcomposer2_4khd, ) has revealed promising capabilities in geometric reasoning, demonstrating competence in solving elementary geometric problems. Parallel developments in external formalized expert systems (unigeo_geoformer/chen2022unigeo, ; unimath/liang2023unimath, ; pgpsnet/zhang2023multi, ; inter-gps/lu2021inter, ) have extended problem solving capabilities to more complex geometric configurations, with highly specialized models (alphageometry/Trinh2024, ; alphageometry2/chervonyi2025, ; tonggeometry/Zhang2024ProposingAS, ) even approaching challenges of the international mathematical olympiad (IMO) level. However, these achievements are critically dependent on high-quality geometric data, currently in short supply. The existing datasets are mainly derived from manual annotations of secondary school textbook problems (geoqa_ngs/chen2021geoqa, ; unigeo_geoformer/chen2022unigeo, ; pgdp5k/hao2022pgdp5k, ; pgpsnet/zhang2023multi, ; inter-gps/lu2021inter, ), and this process is limited in scale and diversity. Although recent integration of MLLMs (e.g. GPT series (gpt4/achiam2023gpt, ; openai2023gpt4v, ; openai2024gpt4o, )) for synthetic question-answer generation (mavis/zhang2024mavis, ; g-llava/gao2023g, ) has partially ameliorated data scarcity concerns, significant challenges persist in ensuring stepwise solution validity [^2] Specifically, automatically generated procedural labels frequently contain latent inconsistencies, such as misapplied geometric axioms or diagram-coordinate mismatches., which may undermine the logical coherence of reasoning chains during model training.
TABLE I: Comparison of various geometric problem-solving datasets and benchmarks. Modality abbreviations: I for Image, FS for Formal Solution, NS for Natural Solution, A for Answer. Thinking Template abbreviations: D for Deductive, BT for Backtrack, MS for Multi-solution.
Datasets & Benchmarks Size Modality Level Thinking Template Labeling Method
GeoQA geoqa_ngs/chen2021geoqa 4998 I, NS, A Middle School D Human Annotation
GeoQA+ dpe-gps/cao2022augmented 7528 I, NS, A Middle School D Human Annotation
Geometry3K inter-gps/lu2021inter 3002 I, FS, A Middle School D Step-wise Formal Verified
PGPS9K pgpsnet/zhang2023multi 9022 I, FS, A Middle School D Step-wise Formal Verified
UniGeo unigeo_geoformer/chen2022unigeo 9543 I, NS, A Middle School D From Source File
GeoEval zhang2024geoeval 5050 I, A Middle & High School - LLM Labeling
MathVista mathvista/lumathvista 1320 * I, A Middle School - From Source File
MathVerse zhang2024mathverse 1745 * I, NS, A High School - LLM Labeling
GeoSense leonardi2024geosence 1789 I, NS, A Synthetic D LLM Labeling & Human Annotation
OlympiadBench he2024olympiadbench 8476 I, NS, A Olympiad-level D From Source File
MAVIS mavis/zhang2024mavis 588K I, NS, A Synthetic D LLM Labeling
G-LLAVA g-llava/gao2023g 170K I, NS, A Synthetic D LLM Lableling
TrustGeoGen 200K I, FS, NS, A Synthetic D, MS, BT Step-wise Formal Verified
GeoTrust-train 2342 I, FS, NS, A Synthetic D, MS, BT Step-wise Formal Verified
GeoTrust-test 240 I, FS, NS, A Synthetic D Step-wise Formal Verified
* These values are not the total count but the count of the GPS category within the dataset.
However, as depicted in table˜I, current geometric datasets exhibit four deficiencies: 1) Modality Fragmentation: The existing datasets suffer from incomplete modality alignment, lacking annotations that synchronize images, textual captions, questions, and chain-of-thought reasoning processes within unified instances. 2) Scalability Issues: Most approaches rely on repurposing human-authored problems from textbooks or examinations rather than generating novel, curriculum-aligned challenges. 3) Cost Constraints: Manual annotation requires costly geometry expertise, while generating solutions through MLLMs consumes excessive computational resources. 4) Trust Deficits: Model-generated solutions lack reliable automated verification mechanisms, introducing risks of undetected logical errors or misaligned reasoning steps.
To this end, we develop TrustGeoGen, a formal language-verified geometric data generation engine that autonomously produces geometric problems, diagrams, and stepwise solutions through four integrated components: 1) a Constructor to generate constraint-satisfying problem premises and diagrams, 2) a Reasoner expanding geometrically valid reasoning graphs with rigorous node verification, 3) a Sampler leveraging the algorithm GeoExplore to extract high-quality reasoning paths, which are later converted into problems by pairing initial premises with derived conclusions, and 4) A Translator converting the data constructed by the formal engine into natural language, and establishing "connection thinking" to bridge the gap between formal reasoning steps and natural human logic. Furthermore, we improve the problem complexity through a bootstrapping mechanism in TrustGeoGen, where formally verified geometric premises in prior stages are recursively fed back to the Constructor to iteratively append new premises. This closed-loop process enables systematic escalation of difficulties while maintaining logical consistency. Additionally, we refine the algorithm GeoExplore to simultaneously explore alternative solution paths and self-correcting reasoning trajectories, generating problems with multiple valid solutions and self-reflective traceback data.
Our proposed TrustGeoGen ensures: 1) Modality Completeness with synchronized dual-form (formal & informal) captions, questions, solutions, and diagrams, 2) Scalability in generating geometry problems by bootstrapping mechanisms without reliance on existing data source and 3) Verifiability through formal validation of all solution steps in its integrated reasoner. Leveraging TrustGeoGen, we synthesized a raw dataset of 200K problems, from which we curated the GeoTrust-train set (2,342 samples) and the GeoTrust-test set (240 samples).
@@
Our experimental results demonstrate that the data generated by TrustGeoGen not only reaches a difficulty level comparable to the international mathematical olympiad (IMO) but also contains novel problems unseen in the pre-training corpora of state-of-the-art (SOTA) models, thereby enhancing their performance. Furthermore, the "connection thinking" provided by the translator bridges the gap between formal engine-verified reasoning paths and human-like logical inference, significantly boosting the model’s geometric problem-solving abilities. The data constructed from "thinking templates" by the GeoExplore series of algorithms has also been proven effective in improving the model’s reasoning capabilities. Additionally, our training data from GeoTrust enhances model performance on the out-of-distribution (OOD) testset GeoQA (geoqa_ngs/chen2021geoqa, ), Geometry3K inter-gps/lu2021inter , and OlympiadBench he2024olympiadbench , highlighting its broader generalization capability. The main contributions of this paper can be summarized as follow:
@@
  • We develop TrustGeoGen, a formalized geometric data generation engine, producing multimodal-integrated data (geometric diagrams, formal & natural language captions, questions and solutions) with trustworthy reasoning guarantees via formal verification.
  • We introduce "Connection Thinking" to bridge the gap between formal and human-like logical reasoning, enabling our engine to enhance models’ reasoning capabilities. Our GeoExplore series algorithms further synthesize data with diverse thinking templates, offering different problem-solving pathways.
  • Extensive experiments show that TrustGeoGen generates challenging, verifiable data that significantly improves models’ geometric problem-solving abilities. Results on OOD datasets demonstrate the effectiveness and generalization of our synthetic data.
2 Related Work
2.1 Advances in Multimodal Large Language Model
Recent advances in large language models (LLMs) have demonstrated remarkable progress in linguistic intelligence, achieving human-level capabilities in various applications (chatgpt/ouyang2022training, ; llama/touvron2023llama, ; llama2/touvron2023llama, ; internlm/team2023internlm, ). Building upon this foundation, the research community has focused on extending these text-based architectures to process visual information, leading to the emergence of sophisticated multimodal large language models (MLLMs) (qwen-vl/bai2023qwen, ; gpt-4v/achiam2023gpt, ; reid2024gemini, ). These integrated systems typically employ modality alignment mechanisms, such as Q-former (blip-2/li2023blip, ) and linear layers (llava/liu2024visual, ), to establish connections between visual representations and text embeddings. Recently, an important application of MLLMs is their ability to be regarded as multimodal agents that interact with scientific environments (team2025novelseek, ; zhang2025origene, ; gottweis2025towards, ), thereby generating valuable scientific discoveries and driving productivity advancements. Although current MLLMs show strong performance in standard vision-language benchmarks (fu2023mme, ; xia2024docgenome, ; structchart/xia2023structchart, ; mathvista/lumathvista, ; jiang2025mme, ), their effectiveness decreases significantly when processing mathematical visualizations that require reasoning. Recent specialized approaches address this limitation through targeted training strategies. For example, geometric reasoning capabilities have been enhanced through domain-specific datasets containing annotated diagrams (g-llava/gao2023g, ; mavis/zhang2024mavis, ).
2.2 Geometric Problem Solving
As a challenging task, geometric problem solving requires understanding diagrams, interpreting symbols, and performing complex reasoning. While MLLMs have shown remarkable proficiency in human-level reasoning and generation capabilities, they struggle with automatic geometric problem solving due to their pre-training on natural images and texts, as well as the lack of automated verification in the problem-solving process. To address this limitation, researchers have proposed various approaches to improve the understanding of geometric images and corpora, including unimodal pre-training, vision-language alignment, and visual instruction tuning (geox/Xia2024GeoXGP, ; mavis/zhang2024mavis, ; g-llava/gao2023g, ; geoqa_ngs/chen2021geoqa, ; unigeo_geoformer/chen2022unigeo, ; pgpsnet/zhang2023multi, ; pgdp5k/hao2022pgdp5k, ; jiang2025mme, ). Another major approach involves the use of formalized solvers to tackle geometric problems. While these solvers (alphageometry/Trinh2024, ; alphageometry2/chervonyi2025, ; sicca2024newclid, ) are capable of addressing challenges at the level of the international mathematical olympiad (IMO), they require precise modeling of geometric problems to fit the solver’s language. This necessity introduces significant obstacles in terms of practical application and generalization, limiting the usability in real-world scenarios.
2.3 Datasets and Benchmarks
High-quality training data is crucial for enhancing the geometry problem-solving capabilities of MLLMs. As summarized in table˜I, there are three common data construction approaches. The first approach involves filtering real-world data and then annotating it manually geoqa_ngs/chen2021geoqa ; dpe-gps/cao2022augmented ; unigeo_geoformer/chen2022unigeo ; mathvista/lumathvista ; he2024olympiadbench . While this method yields high-quality multi-modal reasoning data, its scalability is limited by the finite data pool and the intensive labor required. Furthermore, as problem complexity increases, the difficulty of annotation rises sharply, typically constraining the resulting problems to a lower difficulty range. The second approach utilizes formal engines for data generation inter-gps/lu2021inter ; pgpsnet/zhang2023multi . This allows for the rapid creation of large-scale datasets with verifiably correct reasoning processes. However, the quality of the data is inherently tied to the sophistication of the formal engine. A significant gap often exists between the synthesized logic and authentic human reasoning, leaving considerable room for improvement in enhancing the model’s natural language inference abilities. Finally, the third method employs LLMs to synthesize reasoning data mavis/zhang2024mavis ; g-llava/gao2023g . This method is highly efficient, and the generated reasoning steps more closely resemble natural language. Its primary drawback, however, is that the outputs are unverifiable; they may contain erroneous intermediate steps that could mislead or corrupt the model’s reasoning capabilities.
@@ 3.1 Data Engine
In the process of generating multi-modal geometric reasoning data, it employs a rule-based data construction method augmented by a geometric verifier to synthesize high-quality datasets. This approach ensures that the generated data are accurate and controllable. As shown in  fig.˜1, our framework consists of four components: Constructor, Reasoner, Sampler, and Translator.
Constructor is responsible for generating a geometric scene for a given problem. The foundational components of this process are constructions ($C$) and statements ($s$). A geometric scene is described by a sequence of constructions, where each construction $C_{i}$ generates one or more statements, such as $\{s_{i}^{1},s_{i}^{2},$...\}, that define the relationships between geometric elements. Before a new construction can be added to the scene, the existing statements must satisfy its preconditions. For instance, to apply the construction "draw a perpendicular from point $A$ to line $BC$, with $D$ as the foot," the precondition "points $A$, $B$, and $C$ are not collinear" must hold true. Conversely, the successful application of a construction yields new statements. For example, the construction "create an isosceles triangle ABCABC" introduces the statements "AB=ACAB=AC" and $\angle ABC=\angle ACB$". Following this paradigm, a complete geometric scene, described by $\{C_{1},C_{2},...,C_{n}\}$, can be built incrementally from scratch. This process concurrently generates the scene’s definitive set of initial statements, $S_{0}=\{s_{1}^{1},s_{1}^{2},...,s_{n}^{1},s_{n}^{2},$...\}, which is the total collection of all generated statements.
To ensure the generated scenes are both geometrically significant and numerically instantiated, we have hand-crafted a library of 46 base scene generation functions. These functions produce primitive geometric configurations populated with random numerical values. As shown in fig.˜1, the Constructor initiates the process by randomly selecting one of these functions to create an initial, numerically-defined scene. Subsequently, it iteratively adds further constructions, contingent on their preconditions being met, until a target scene is fully formed. This scene is then passed to a geometric compiler for validation, which checks for the consistency of all geometric relations and constraints. Ultimately, once a valid geometric scene is successfully constructed, its complete set of initial statements, $S_{0}$, can be obtained.
1
2
Require: Reasoning graph $\mathcal{G}=(S,S_{0},R,\hookrightarrow)$, target statement $s_{n}$, threshold $\tau_{l}$, threshold $\tau_{r}$
Ensure: Filtered reasoning path $\mathcal{P}$
3
4Initialize $\mathcal{P}\leftarrow\emptyset$;
5 $s\leftarrow s_{n}$;
6
7while $s\notin S_{0}$ do
8   Identify rule $r_{s}\in R$ and parent statement set $S_{i-1}\subset S$ such that $S_{i-1}\xhookrightarrow{r_{s}}s$;
9   Append transition $(S_{i-1},r_{s},s)$ to $\mathcal{P}$;
10 $s\leftarrow S_{i-1}$;
11
12
13Assess path length: $L\leftarrow|\mathcal{P}|$;
14 Compute premises ratio: $R_{P}\leftarrow|S_{P}|/|S_{0}|$, where $S_{P}\subseteq\mathcal{P}$;
15
16if $L\geq\tau_{l}$ and $R_{P}\geq\tau_{r}$ then
17 return $\mathcal{P}$;
18
19else
20 Discard $\mathcal{P}$;
21